Home > News > Archive > s20180305
Storm aftermath update 05 March 2018

11th March: The LT is back on-sky after the successful change of the hydraulic oil that runs the telescope's bearings.

The ice that encased the LT and other telescopes since late January appears to have finally melted. The LT's enclosure was tested this morning and found to be in good working order. However we've found water in the telescope's hydraulic oil bearing system which needs to be removed, so the observatory currently remains off-sky for a few days until this can be done.

The first image below shows the enclosure being tested this morning, as taken by our external webcam. The other images were taken by our engineering and maintenance support contractor Dirk Raback over the last few weeks on the occaisons he was able to access the site. They show some of the ice that caused us problems and some of the damage caused by last week's storm. The latest images are shown here, but see our Flickr "Telescope and Site" album for all of Dirk's recent ice and storm images.

5th March: Successful test of the enclosure.

2nd March: Rockfalls on access road to mountaintop. © 2018 Dirk Raback

21st Feb: Ice on enclosure ram mechanisms. © 2018 Dirk Raback

16th Feb: General ice coverage in vicinity of telescope. © 2018 Dirk Raback

12th Feb: Ice on signpost.
© 2018 Dirk Raback